### v1

# NeuroStrata: Harnessing Neurosymbolic Paradigms for Improved Design, Testability, and Verifiability of Autonomous CPS

Abstract.
Autonomous cyber-physical systems (CPSs) leverage AI for perception, planning, and control but face trust and safety certification challenges due to inherent uncertainties. The neurosymbolic paradigm replaces stochastic layers with interpretable symbolic AI, enabling determinism. While promising, challenges like multisensor fusion, adaptability, and verification remain. This paper introduces NeuroStrata, a neurosymbolic framework to enhance the testing and verification of autonomous CPS. We outline its key components, present early results, and detail future plans.
AI-based Systems, Cyber-Physical Systems, Neurosymbolic AI, Testing, Verification
Ivan Ruchkin is supported in part by the NSF Grant CCF-2403616. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation (NSF) or the United States Government.
copyright: rightsretainedccs: Software and its engineering Software verification and validationccs: Computer systems organization Embedded and cyber-physical systemsccs: Computing methodologies Machine learningccs: Theory of computation Formal languages and automata theory
1. CONTEXT, MOTIVATION, AND AIMS
The integration of machine learning (ML) into CPS has driven innovations in autonomous vehicles (Waymo, 2023; Tesla Autopilot, 2023; Uber ATG, 2023), delivery drones (Amazon Prime Air, 2023; Google Wing, 2023; Zipline, 2023), and robotic surgeries (Davinci, 2023; Mazor, 2023; Mako, 2023). While ML enhances autonomy and intelligence, its uncertain and brittle nature, as witnessed in softmax-based classifications and regression-based control, undermines traditional formal verification, necessitating novel solutions.
Figure 1. Motivating system: an industrial-strength software stack for autonomous driving (Deng et al., 2022b).
Neurosymbolic approaches combine symbolic reasoning with neural learning, addressing challenges in autonomous CPSs by enabling co-training of neural components with symbolic logic through probabilistic logic programming and differentiable reasoning (Cohen et al., 2017; Manhaeve et al., 2021; Li et al., 2023; Lu et al., 2024). Advanced methods, such as program synthesis using Domain-Specific Languages (DSLs), demonstrate promise by supporting both deterministic and probabilistic programs (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021). However, significant bottlenecks persist, including unseen data in real-world deployments (Fredrikson et al., 2023), multi-sensor fusion challenges (Strickland et al., 2018), and the lack of neural component verification. Determinism at the decision level of perception and mission planning is also missing, which can enable the applicability of decades-long formal verification and testing techniques. Ensuring system-level safety and liveness requires systematic co-design of perception, planning, and control — a critical aspect missing in current methods.
This paper summarizes state-of-the-art neurosymbolic paradigms, using autonomous driving as a case study to highlight gaps in adaptability to unseen environments, multi-sensor complexity, systematic validation of neural components, and determinism at the decision level, critical for reliable autonomous CPS.
To address these issues, we put forward NeuroStrata — a neurosymbolic framework for designing and assuring autonomous CPSs. Our vision integrates neurosymbolic distillation and corner-case test generation using LLMs to enable data-driven specification mining, top-down synthesis of symbolic and neurosymbolic components, and runtime bottom-up adaptation via program induction. This approach evolves symbolic programs dynamically for decision-making in perception and planning/control modules. Building on prior work, we aim to transform the testability and verifiability of autonomous CPSs through the neurosymbolic paradigm.
2. Motivating System and the State-of-the-art
State-of-the-art autonomous CPSs, such as autonomous driving systems (ADS), are typically built on middleware frameworks like Robot Operating System (ROS) with proprietary extensions (e.g., Baidu’s CyberRT) to reduce message latency (Zheng et al., 2024; Liang et al., 2023; Lou et al., 2022; Deng et al., 2021; Prakash et al., 2021). As shown in Figure 1, these systems integrate perception, prediction, planning, and control modules. The perception module processes multi-modal sensor data (e.g., LiDAR, cameras, IMU) for obstacle detection, traffic light recognition, and localization. The prediction module forecasts dynamic object trajectories, while the planning module computes the ego vehicle’s trajectory. The control module generates actuation commands, such as steering and braking. AI components are pervasive across these modules, enabling object detection, trajectory prediction, and real-time control.
Recent advancements in verifying machine learning models and testing learning-enabled CPS have utilized methods like NNV star sets (Tran et al., 2020), Sherlock (Dutta et al., 2019), Reluplex (Katz et al., 2017), and Branch-and-Bound (Bunel et al., 2018). While promising, these methods often fall short in providing comprehensive safety and liveness guarantees, particularly when dealing with the complexity and scale of the real-world, multi-modal CPS as shown in the motivating system above. System-level testing (Deng et al., 2022a; Tian et al., 2022; Li et al., 2020) and robustness analyses (Tian et al., 2018; Deng et al., 2020; Cai and Koutsoukos, 2020) offer progress but still rely on probabilistic methods to assure probabilistic systems (Cleaveland et al., 2022; Badithela et al., 2023), which can weaken the guarantees (Badings et al., 2023).
In parallel, neurosymbolic paradigms aim to combine neural and symbolic approaches — either through differentiable logic programming to produce differentiable yet deterministic outputs (Li et al., 2023; Raghothaman et al., [n. d.]) or through program induction to synthesize deterministic programs by observing limited input-output pairs (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021). However, they face notable limitations. The differentiable logic programming paradigm struggles with hard-coded mappings between neural component outputs and logic program inputs, as well as the complexity of implementing differentiable logic programming capabilities. These constraints limit its ability to support multi-modal sensors and the sophisticated logics required for autonomous systems, which demand more automated and expressive guidance. Similarly, the program induction paradigm relies on Domain-Specific Languages (DSLs) for guidance, which are often hard-coded and lack the flexibility needed for multi-modal sensor integration and advanced reasoning (Parisotto et al., 2016; Ellis et al., 2021; Bjørner et al., 2023). Furthermore, program induction may not be sufficient to replace all pure AI components in such systems.
Recent work on multimodal neurosymbolic systems integrates visual and auditory signals but relies on simplistic symbolic rules, unsuitable for the diverse sensor modalities of autonomous CPS, such as LiDARs, radars, and cameras (Haq et al., 2024). A Neural State Machine (NSM) approach combines vision and language reasoning but suffers from scalability issues, manual scene graph construction, and limited interpretability due to missing source code (Jain et al., 2024). These limitations underscore the need for refined co-design to meet the stringent requirements of autonomous CPS.
The above limitations emphasize the pressing need for deterministic testing and verification approaches in autonomous CPS, particularly in safety-critical domains like autonomous driving. For instance, perception modules in autonomous vehicles, which rely on uncertain or stochastic processes such as (Bayesian) neural networks for object detection, often fail in “long tail” scenarios where the inputs deviate from training data (covariate shift). Predictable and deterministic approaches, incorporating reasoning layers, can adapt to such unseen scenarios by leveraging symbolic logic to ensure robust decision-making and mitigate failures caused by stochastic uncertainties. This adaptability is crucial for guaranteeing safety and reliability across perception, prediction, and planning modules in dynamic and complex real-world environments (Shalev-Shwartz et al., 2017; Seshia, 2017; Zapridou et al., 2020).
3. NeuroStrata: Our Vision for Hierarchical Neurosymbolic Framework for Autonomous Systems
To address the challenges of designing, testing, and verifying autonomous CPS, we propose a new neurosymbolic framework, NeuroStrata, tailored to the unique requirements of such systems. As shown in Figure 2, NeuroStrata combines neural adaptability with symbolic reasoning to enforce formal specifications across hierarchical DSLs that capture underlying safety and liveness properties. The framework structures Perception and Planning & Control capabilities into high-level (symbolic-only) and middle- and low-level (neurosymbolic) modules. It ensures runtime reliability and adaptation via a two-phase process: top-down synthesis, propagating symbolic specifications to neurosymbolic modules, and bottom-up adaptation, where neurosymbolic outputs refine symbolic programs.
Figure 2. Proposed Vision for NeuroStrata: Hierarchical Neurosymbolic Programming for Autonomous CPS
Modules. At design time, Specification Mining, built on neurosymbolic distillation (Singireddy et al., 2023; Blazek et al., 2024; Abir, 2024), extracts formal safety and liveness specifications from training datasets. To cover more diverse safety and liveness violations and out-of-distribution scenarios beyond existing training data, we leverage recent work using large language models to analyze multi-modal sensor data (Zheng et al., 2024; Deng et al., 2023), such as front-facing cameras in vehicles, to generate additional real-world crashes and unusual cases from various angles. These specifications are propagated hierarchically across the system. In the perception stack, a high-level Scene Graph encodes semantic relationships and interactions between objects (e.g., “pedestrian crossing road”), represented as differentiable, adaptable programs that can be verified using formal tools like theorem provers. The middle-level Semantic Map encodes spatial and semantic information such as road layouts and drivable areas, ensuring consistency with the scene graph via symbolic rules. The low-level Sensor Fusion and Signal Processing integrates multi-modal sensor data (e.g., LiDAR, cameras, GPS) while enforcing constraints on accuracy and consistency, leveraging neurosymbolic reasoning for fusion and processing. Similarly, the planning and control stack follows a hierarchical structure. The high-level Global/Mission Planner synthesizes deterministic programs to achieve overall system objectives, verified with formal methods such as theorem proving. The middle-level Local Planner generates short-term trajectories that align with global plans while adapting to local changes, guided by symbolic reasoning. The low-level Actuation Control converts trajectories into control commands (e.g., steering angle, throttle) and ensures compliance with constraints using runtime verification techniques.
Specifications. For perception, high-level specifications govern system-wide context awareness, such as ensuring that pedestrians and vehicles do not spatially overlap in the scene graph or that all objects adhere to semantic relationships. Middle-level specifications enforce localized consistency, such as aligning lane boundaries with the semantic map and ensuring that detected objects are positioned correctly within the road layout. Low-level specifications address operational constraints, such as maintaining sensor fusion accuracy within a 0.1-meter error margin and ensuring consistent integration of multi-modal sensor data. For planning/control, high-level specifications ensure system-wide safety and mission compliance, such as requiring the vehicle to remain within designated route bounds throughout its journey. Middle-level specifications enforce trajectory-level constraints, such as avoiding obstacles within a 2-meter radius or maintaining smooth transitions between trajectory points. Low-level specifications govern detailed actuation control, such as keeping the steering angle within physical limits and ensuring the stability of throttle and braking in response to control inputs. These hierarchical specifications for perception and planning/control ensure an integrated and reliable system design.
Adaptation. During runtime, NeuroStrata dynamically adapts its perception and planning modules to real-time data while maintaining formal specification compliance. For perception, sensor data flows upward through the hierarchy, where outputs from the low-level sensor fusion are validated against middle-level semantic map constraints, and updates propagate to the high-level scene graph. It evolves dynamically using differentiable program induction, compacting, and adapting specifications as needed. For planning and control, high-level mission planners adjust strategies based on changing conditions, while differentiable and adaptable control programs refine global plans and compact themselves in response to system data. Middle- and low-level components, such as local planners and actuation control, remain guided by symbolic reasoning to ensure safety and alignment with global objectives. This integration enables simultaneously adaptable and formally validated behavior throughout the system.
Guarantees. NeuroStrata ensures reliability through a hybrid validation framework. High-level deterministic programs, such as scene graphs and mission planners, are validated using formal verification tools like model checking and theorem proving. Middle- and low-level neurosymbolic components, such as semantic maps and sensor fusion, are guided by symbolic constraints and validated using white-box testing, runtime monitors, and error propagation analysis (e.g., approximate reachability verification (Geng et al., 2024) and conformance checking (Habeeb et al., 2024)). Together, this framework bridges the gap between deterministic high-level programs and adaptive, data-driven neuro-components, thus providing formal guarantees across all three levels of the hierarchy.
4. EARLY RESULTS AND FUTURE PLAN
We conducted a preliminary case study to investigate a key Research Question (RQ): “can neurosymbolic reasoning complement neural-network training to align with underlying specifications?”. We also outline our future plans, along with the potential challenges and proposed solutions.
4.1. Assessing neurosymbolic reasoning to align neural network training with specifications
In this study, we investigate the capability of differentiable neurosymbolic reasoning to align perceptual neural networks with specifications. We explore the application of a high-level visual perception system trained by aligning its output with specifications. In this application, the goal is to infer spatio-temporal scene graphs (STSG) from videos (e.g., ones taken by ego-centric cameras), where the scene graphs must align with a given spatio-temporal specification. Figure 3 depicts a specification for a traffic scene which is described in natural language but then formalized into a temporal logic formula. Notice that the specification consists of logical symbols like exists ($\exists$), and ($\wedge$), not ($\neg$¬), and finally ($\Diamond$). A neurosymbolic approach to solving this task comprises of a neural model for STSG extraction and a differentiable symbolic component for aligning the predicted STSG with the given specification. Being differentiable, the loss computed from the alignment process can be used to supervise the neural model. We evaluated our work on three datasets: OpenPVSG (Yang et al., 2023), 20BN-Something-Something (Goyal et al., 2017), and MUGEN (Hayes et al., 2022), each with diverse temporal properties. Our approach outperforms current baselines on downstream tasks while offering explainability [^1] Reference to the full report was anonymized for the review process.. This specification alignment provides high confidence in our top-down synthesis approach and in guiding neural training with our specifications.
Figure 3. Aligning STSG with natural language description via temporal logic specifications.
4.2. Future Research Plan
To advance NeuroStrata, we propose a six-step future plan with concrete steps to address challenges at each stage.
First, generating diverse training datasets will leverage recent advancements in model-based testing that utilize LLMs to analyze multi-modal sensor data (Zheng et al., 2024; Deng et al., 2023). The multi-faceted challenge lies in ensuring the generated datasets are diverse, representative of real-world scenarios, and capable of addressing edge cases. Solutions can be tailored around prior work by accessing diverse sensor datasets and logs, leveraging advanced multi-modal LLMs, and integrating domain-specific constraints with iterative refinement based on industrial partner feedback.
Second, designing suitable DSLs is essential for capturing hierarchical and semantically rich specifications. These DSLs enable experts to encode operational constraints for sensor fusion, signal processing, and physical actuation control. Challenges include ensuring the DSLs are intuitive for domain experts while expressive enough to handle complex requirements. Solutions involve co-designing DSLs with autonomy and robotics specialists, developing language automation, and designing usable visual interfaces. By providing a bridge between formal methods and practical application, these DSLs empower experts to play an active role in system design.
Third, developing a specification mining module based on neurosymbolic distillation will extract formal safety and liveness specifications from training datasets and LLM interactions. Aligning mined specifications with real-world requirements and handling noisy/hallucinated data are key challenges. Hybrid approaches that combine symbolic reasoning with neural embeddings, as well as active learning techniques, can iteratively refine the mined specifications to ensure accuracy and physical grounding.
Fourth, for design-time synthesis and verification, we will enforce multi-level specifications for perception and planning/control modules, leveraging the hierarchical structure defined in our DSL. Modular architectures will enable scalable top-down synthesis of symbolic and neurosymbolic components, while formal verification ensures compliance with specifications. Parallelized processes and adaptive abstraction techniques will address scalability challenges, ensuring robustness across diverse scenarios and high-dimensional inputs.
Fifth, for runtime adaptation and validation, we will develop mechanisms to dynamically refine symbolic programs for real-world changes while ensuring specification compliance. Inspired by program induction approaches like DreamCoder (Ellis et al., 2021), NeuroStrata will iteratively refine symbolic representations using real-time data. Challenges include maintaining computational efficiency and real-time guarantees. To address these, we will optimize runtime validators, integrate lightweight symbolic reasoning for faster adaptation, and implement efficient runtime verification to ensure reliability and compliance with minimal overhead. These advancements will enable NeuroStrata to adapt to dynamic environments and evolving operational conditions.
Finally, for industrial deployment, NeuroStrata will be applied to autonomous driving systems, delivery drones, cargo drones, and passenger aircraft — as facilitated by our partners. Key challenges include seamless integration into existing systems, adherence to stringent safety standards, and building trust among stakeholders. Solutions include close collaborative projects, iterative deployment in increasingly open environments, and the creation of comprehensive documentation and training programs to facilitate adoption.
5. CONCLUSION
This paper explores the potential and challenges of neurosymbolic paradigms for designing, testing, and verifying autonomous CPS. We propose NeuroStrata, a framework enabling top-down synthesis of symbolic and neurosymbolic components for perception and planning/control, and bottom-up adaptation of symbolic programs for real-time decisions. Early results validate neural alignment with specifications. We outline challenges and solutions for implementing NeuroStrata, aiming to bridge theoretical advancements and practical applications, transforming autonomous CPS testing and verification in real-world scenarios.