### v2

# NeuroStrata: Harnessing Neurosymbolic Paradigms for Improved Design, Testability, and Verifiability of Autonomous CPS

Abstract.
Autonomous cyber-physical systems (CPSs) leverage AI for perception, planning, and control but face trust and safety certification challenges due to inherent uncertainties. The neurosymbolic paradigm replaces stochastic layers with interpretable symbolic AI, enabling determinism. While promising, challenges like multisensor fusion, adaptability, and verification remain. This paper introduces NeuroStrata, a neurosymbolic framework to enhance the testing and verification of autonomous CPS. We outline its key components, present early results, and detail future plans.
AI-based Systems, Cyber-Physical Systems, Neurosymbolic AI, Testing, Verification
This work is supported in part by the Australian Research Council (FT240100269, DP210102447, LP190100676), the U.S. National Science Foundation (CCF 2403616, NAIAD 2332744), and the National AI Institute for Edge Computing Leveraging Next Generation Wireless Networks (CNS 2112562).
journalyear: 2025copyright: ccconference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728500isbn: 979-8-4007-1276-0/2025/06ccs: Software and its engineering Software verification and validationccs: Computer systems organization Embedded and cyber-physical systemsccs: Computing methodologies Machine learningccs: Theory of computation Formal languages and automata theory
1. CONTEXT, MOTIVATION, AND AIMS
The integration of machine learning (ML) into cyber-physical systems (CPS) has spurred innovations in autonomous vehicles (Waymo, 2023; Tesla Autopilot, 2023; Uber ATG, 2023), delivery drones (Amazon Prime Air, 2023; Google Wing, 2023; Zipline, 2023), and robotic surgery (Davinci, 2023; Mazor, 2023; Mako, 2023). While ML enhances autonomy, its inherent uncertainty and brittleness—evident in softmax-based classifications and regression-based control—limit the applicability of traditional formal verification, calling for new approaches. Neurosymbolic methods address this gap by combining neural learning with symbolic reasoning, enabling co-training via probabilistic logic programming and differentiable reasoning (Cohen et al., 2017; Manhaeve et al., 2021; Li et al., 2023; Lu et al., 2024). Recent advances, including DSL-based program synthesis, support both deterministic and probabilistic behaviors (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021).
Despite progress, key challenges remain: handling unseen deployment data (Fredrikson et al., 2023), multi-sensor fusion (Strickland et al., 2018), and the absence of neural component verification and decision-level determinism. These limitations hinder the application of mature formal methods to autonomous CPS. Ensuring safety and liveness requires systematic co-design of perception, planning, and control—still lacking in current frameworks.
This paper reviews state-of-the-art neurosymbolic paradigms, using autonomous driving to illustrate limitations in adaptability, sensor integration, component validation, and decision determinism—each critical for reliable CPS operation. To address these, we propose NeuroStrata—a neurosymbolic framework for the design and assurance of autonomous CPSs. It integrates neurosymbolic distillation and LLM-guided test generation for data-driven specification mining, top-down synthesis of hybrid components, and runtime bottom-up adaptation via program induction. This enables dynamic evolution of symbolic programs within perception and control modules, advancing the testability and verifiability of autonomous CPSs.
Figure 1. Motivating system: an industrial-strength software stack for autonomous driving (Deng et al., 2022b).
2. Motivating System and State-of-the-art
State-of-the-art autonomous CPSs, such as autonomous driving systems (ADS), are typically built on middleware frameworks like Robot Operating System (ROS) with proprietary extensions (e.g., Baidu’s CyberRT) to reduce message latency (Zheng et al., 2024; Liang et al., 2023; Lou et al., 2022; Deng et al., 2021; Prakash et al., 2021). As shown in Figure 1, these systems integrate perception, prediction, planning, and control modules. The perception module processes multi-modal sensor data (e.g., LiDAR, cameras, IMU) for obstacle detection, traffic light recognition, and localization. The prediction module forecasts dynamic object trajectories, while the planning module computes the ego vehicle’s trajectory. The control module generates actuation commands, such as steering and braking. AI components are pervasive across these modules, enabling object detection, trajectory prediction, and real-time control.
Recent tools for verifying neural networks, such as NNV star sets (Tran et al., 2020), Sherlock (Dutta et al., 2019), and Reluplex (Katz et al., 2017), focus primarily on verifying local and input-output properties, which are too narrow to capture system-level safety. Some efforts build on these tools for system-level safety (Ivanov et al., 2021; Geng et al., 2024) but are far from verifying real-world systems. While system-level testing (Deng et al., 2022a; Tian et al., 2022; Li et al., 2020) and robustness analysis (Tian et al., 2018; Deng et al., 2020; Cai and Koutsoukos, 2020) offer broader insights, they often rely on probabilistic reasoning (Cleaveland et al., 2022; Badithela et al., 2023), which weakens formal guarantees (Badings et al., 2023).
In parallel, neurosymbolic paradigms aim to unify neural and symbolic methods. One way to do so is via differentiable logic programming for producing deterministic yet trainable outputs (Li et al., 2023; Raghothaman et al., [n. d.]). Another approach is program induction that synthesizes symbolic programs from limited input-output examples (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021).
However, neurosymbolic approaches face non-trivial challenges. Differentiable logic programming often requires hard-coded mappings between neural outputs and logical inputs, and its implementation complexity restricts sensor multi-modality and expressive logics needed in autonomous systems. Likewise, program induction relies on fixed domain-specific languages (DSLs) that lack flexibility for sensor fusion or advanced reasoning (Parisotto et al., 2016; Ellis et al., 2021; Bjørner et al., 2023). Our vision aligns with neurosymbolic reinforcement learning (NS-RL), which combines deep learning with symbolic reasoning for control (Acharya et al., 2023); however, NeuroStrata proposes a broader architecture-level redesign. It supports formal specification, verification, and adaptation across perception, planning, and control, with NS-RL serving as a possible component at lower levels. NeuroStrata enables end-to-end reasoning, design-time synthesis, runtime validation, and system-level verification over properties expressed in rich logics.
Neurosymbolic distillation is increasingly adopted to unify neural learning with symbolic interpretability. For example, Automaton Distillation (Singireddy et al., 2023) maps Q-values from deep RL agents into deterministic finite-state automata for policy transfer, while NUDGE (Delfosse et al., 2023) distills neural policies into weighted first-order logic rules using neurally guided abstraction and differentiable forward reasoning. In contrast, NeuroStrata targets broader system-level specifications across perception, planning, and control, and uniquely integrates such distillation with large language models and active learning via L* to extract and refine signal temporal logic specifications, as demonstrated in our prior work (Zheng et al., 2024).
Recent work on multimodal neurosymbolic systems integrates visual and auditory signals but relies on simplistic symbolic rules, unsuitable for the diverse sensor modalities of autonomous CPS, such as LiDARs, radars, and cameras (Haq et al., 2024). A Neural State Machine (NSM) approach combines vision and language reasoning but suffers from scalability issues, manual scene graph construction, and limited interpretability due to missing source code (Jain et al., 2024). These limitations underscore the need for refined co-design to meet the stringent requirements of autonomous CPS.
The above limitations emphasize the pressing need for deterministic testing and verification approaches in autonomous CPS, particularly in safety-critical domains like autonomous driving. For instance, perception modules in autonomous vehicles, which rely on uncertain or stochastic processes such as (Bayesian) neural networks for object detection, often fail in “long tail” scenarios where the inputs deviate from training data (covariate shift). Predictable and deterministic approaches, incorporating reasoning layers, can adapt to such unseen scenarios by leveraging symbolic logic to ensure robust decision-making and mitigate failures caused by stochastic uncertainties. This adaptability is crucial for guaranteeing safety and reliability across perception, prediction, and planning modules in dynamic and complex real-world environments (Shalev-Shwartz et al., 2017; Seshia, 2017; Zapridou et al., 2020).
3. NeuroStrata: Our Vision for Hierarchical Neurosymbolic Framework for Autonomous Systems
To address the challenges of designing, testing, and verifying autonomous CPS, we propose a new neurosymbolic framework, NeuroStrata, tailored to the unique requirements of such systems. As shown in Figure 2, NeuroStrata combines neural adaptability with symbolic reasoning to enforce formal specifications across hierarchical DSLs that capture underlying safety and liveness properties. The framework structures Perception and Planning & Control capabilities into high-level (symbolic-only) and middle- and low-level (neurosymbolic) modules. It ensures runtime reliability and adaptation via a two-phase process: top-down synthesis, propagating symbolic specifications to neurosymbolic modules, and bottom-up adaptation, where neurosymbolic outputs refine symbolic programs.
Figure 2. Proposed Vision for NeuroStrata: Hierarchical Neurosymbolic Programming for Autonomous CPS
Modules. At design time, Specification Mining, built on neurosymbolic distillation (Singireddy et al., 2023; Blazek et al., 2024; Delfosse et al., 2023), extracts formal safety and liveness specifications from training datasets. To cover more diverse safety and liveness violations and out-of-distribution scenarios beyond existing training data, we leverage recent work using large language models to analyze multi-modal sensor data (Zheng et al., 2024; Deng et al., 2023), such as front-facing cameras in vehicles, to generate additional real-world crashes and unusual cases from various angles. These specifications are propagated hierarchically across the system. In the perception stack, a high-level Scene Graph encodes semantic relationships and interactions between objects (e.g., “pedestrian crossing road”), represented as differentiable, adaptable programs that can be verified using formal tools like theorem provers. The middle-level Semantic Map encodes spatial and semantic information such as road layouts and drivable areas, ensuring consistency with the scene graph via symbolic rules. The low-level Sensor Fusion and Signal Processing integrates multi-modal sensor data (e.g., LiDAR, cameras, GPS) while enforcing constraints on accuracy and consistency, leveraging neurosymbolic reasoning for fusion and processing. Similarly, the planning and control stack follows a hierarchical structure. The high-level Global/Mission Planner synthesizes deterministic programs to achieve overall system objectives, verified with formal methods such as theorem proving. The middle-level Local Planner generates short-term trajectories that align with global plans while adapting to local changes, guided by symbolic reasoning. The low-level Actuation Control converts trajectories into control commands (e.g., steering angle, throttle) and ensures compliance with constraints using runtime verification techniques.
Specifications. For perception, high-level specifications govern system-wide context awareness, such as ensuring that pedestrians and vehicles do not spatially overlap in the scene graph or that all objects adhere to semantic relationships. Middle-level specifications enforce localized consistency, such as aligning lane boundaries with the semantic map and ensuring that detected objects are positioned correctly within the road layout. Low-level specifications address operational constraints, such as maintaining sensor fusion accuracy within a 0.1-meter error margin and ensuring consistent integration of multi-modal sensor data. For planning/control, high-level specifications ensure system-wide safety and mission compliance, such as requiring the vehicle to remain within designated route bounds throughout its journey. Middle-level specifications enforce trajectory-level constraints, such as avoiding obstacles within a 2-meter radius or maintaining smooth transitions between trajectory points. Low-level specifications govern detailed actuation control, such as keeping the steering angle within physical limits and ensuring the stability of throttle and braking in response to control inputs. These hierarchical specifications for perception and planning/control ensure an integrated and reliable system design.
Adaptation. At runtime, symbolic reasoning and neural components interact through a bi-directional feedback loop. Neural modules (e.g., sensor fusion and perception) generate scene-level representations, which are continuously validated against symbolic specifications such as semantic constraints and temporal logic. These neural outputs trigger updates to symbolic programs (e.g., the scene graph program and global planning program) via differentiable program induction (see Figure 2). This progressively evolves symbolic components in response to new observations. In turn, symbolic planners guide neural policy adaptation and ensure specification compliance at all layers. This closed neuro-symbolic loop supports runtime behavior that is both adaptive and formally grounded.
Guarantees. NeuroStrata ensures reliability through a hybrid validation framework. High-level deterministic programs, such as scene graphs and mission planners, are validated using formal verification tools like model checking and theorem proving. Middle- and low-level neurosymbolic components, such as semantic maps and sensor fusion, are guided by symbolic constraints and validated using white-box testing, runtime monitors, and error propagation analysis (e.g., approximate reachability verification (Geng et al., 2024) and conformance checking (Habeeb et al., 2024)). Together, this framework bridges the gap between deterministic high-level programs and adaptive, data-driven neuro-components, thus providing formal guarantees across all three levels of the hierarchy.
4. EARLY RESULTS AND FUTURE PLAN
We conducted a preliminary case study to explore the key Research Question (RQ): “Can neurosymbolic reasoning complement neural network training to better align with specifications?” We also outline future directions, challenges, and proposed solutions.
4.1. Specification-Aligned Training
[…]
4.2. Advancing NeuroStrata: A Six-Step Plan
To operationalize NeuroStrata, we propose a six-step roadmap that addresses key challenges in specification-driven learning and adaptation for AI-based CPS.
I. Diverse training dataset generation will leverage model-based scenario synthesis using large language models, as demonstrated in (Zheng et al., 2024; Elmaaroufi et al., 2024), and validated on real systems in (Deng et al., 2023). This approach mitigates covariate shift by generating multi-modal, edge-case–rich training and testing datasets, grounded in physical realism and enhanced with domain-specific constraints.
II. Domain-specific languages (DSLs) will be co-designed with robotics and autonomy experts to express hierarchical specifications across sensing, planning, and control. These DSLs serve as a bridge between formal methods and system engineering, ensuring usability while enabling precise constraint encoding for downstream reasoning and synthesis.
III. Specification mining via neurosymbolic distillation will extract logical properties from training data and LLM queries. To manage hallucinated or noisy outputs, we will integrate symbolic-neural hybrid techniques and active learning to iteratively refine mined specifications for correctness and interpretability.
IV. During design-time synthesis and verification, symbolic and neurosymbolic modules will be synthesized from DSL-defined specifications and verified using compositional and abstraction-based methods. Parallel verification and modular analysis will allow scalability to high-dimensional state spaces, ensuring robust deployment in complex CPS environments.
V. For runtime adaptation and validation, NeuroStrata will employ differentiable program induction (Ellis et al., 2021) to update symbolic programs on-the-fly. To ensure computational feasibility, we will build on the efficient runtime monitoring infrastructure in (Zheng et al., 2016), which uses low-overhead logging and an aspect-oriented programming paradigm to implement scalable runtime verifiers. This setup supports dynamic refinement of neural-symbolic components with minimal latency, addressing both real-time constraints and covariate shift at runtime.
VI. For industrial deployment, NeuroStrata will transition from simulation-based testing in CARLA, and AirSim-based environments to real-world integration with autonomous vehicles, uncrewed aerial vehicles, and robotics platforms. We have initiated research agreements with industry partners in these domains to support this transition. Our upcoming Shonan Seminar (No. 235) and Dagstuhl Seminar (No. 202501048) will further accelerate industry engagement and foster collaborative development and deployment of neurosymbolic CPS solutions.
Figure 3. Aligning STSG with natural language description via temporal logic specifications.
5. CONCLUSION
This paper investigates the potential of neurosymbolic paradigms in the design, testing, and verification of autonomous CPS. We introduce NeuroStrata, a framework that supports top-down synthesis of symbolic and neurosymbolic components for perception and control, and bottom-up adaptation of symbolic programs for real-time decision-making. Preliminary results demonstrate neural alignment with specifications. We discuss key challenges and implementation strategies, aiming to bridge theory and practice in autonomous CPS verification.